# Chromatic numbers, Buchstaber numbers and chordality of Bier spheres

### v1

@@ 3. Bier spheres and their chromatic numbers
In this section, we show that every Bier sphere $\mathrm{Bier}(K)$ of a simplicial complex $K$ on $[m]$ with $m\geq 2$ has the chromatic number equal to either $m-1$, or $m$ and describe all the simplicial complexes $K$, for which $\chi(\mathrm{Bier}(K))=m-1$. The discussion of chromatic numbers of 1-dimensional Bier spheres in the end of the previous section as well as Example 2.11 will serve as the first step in the desired description.
@@
$i,i^{\prime}\mapsto c_{i},\text{ for }1\leq i\leq m,$
@@
Obviously, any cone over a simplicial complex is a weak cone complex, and any suspension over a simplicial complex is a weak suspension complex. The opposite implications fail, as the following example illustrates.
@@
Figure 3. Octahedron is a suspension and therefore a weak suspension, but tetrahedron and icosahedron are not even weak suspensions.
Here is a family of simplicial spheres, one sphere for each $m\geq 5$, such that each member of the family is a weak suspension, but not a suspension.
@@
By definition, its Alexander dual complex $K_{m}^{\vee}$ is a simplicial complex on the vertex set $[m^{\prime}]$ given by
[…]
@@
It is easy to see that $C(m)$ provides a proper coloring of both $K_{m}$ and $K_{m}^{\vee}$ with $(m-1)$ colors, and so
@@
On the other hand, the 1-skeleta of the full subcomplexes $K_{m}\cap 2^{[m-1]}$ and $K^{\vee}_{m}\cap 2^{[(m-1)^{\prime}]}$ are complete graphs with $(m-1)$ vertices, and therefore
@@
Skeleta of a simplex never yield a Bier sphere, which is a suspension, as shown in the following example. In the case of the 0-skeleton of the $m$-simplex and its Alexander dual, which is the $(m-3)$-skeleton of the $m$-simplex, the Bier sphere is not even a weak suspension, since it does not contain any ghost vertices, nor does it contain both edges $\{i,j\}$ and $\{i^{\prime},j^{\prime}\}$, for each pair of distinct vertices $i,j\in[m]$. The next example shows that in all other cases, the Bier sphere of a skeleton of a simplex is a weak suspension and not a suspension.
@@
Let $m\geq 5$, $41\leq k\leq m-$ 4, and $K=\Delta_{[m]}^{(k)}$, i.e., the set of maximal simplices is $\mathrm{M}(K)=\{I\subset[m]\,|\,|I|=k+1\}$. Then $I^{\prime}$ is a minimal nonsimplex in $K^{\vee}$ if and only if $|I|=m-k-2$, hence $K^{\vee}=\Delta_{[m^{\prime}]}^{(m-k-3)}$. For every $i,j\in[m]$, $i\neq j$, we have $\{i,j^{\prime}\}$, $\{i^{\prime},j\}\in\mathrm{Bier}(K)$, since there are no ghost vertices in $K$ and $K^{\vee}$, $\{i,j\}\in K\subset\mathrm{Bier}(K)$, since $k\geq 1$, and $\{i^{\prime},j^{\prime}\}\in K^{\vee}\subset\mathrm{Bier}(K)$, since $m-k-3\geq 1$. Therefore $\mathrm{Bier}(K)$ is a weak suspension for every pair $\{i,i^{\prime}\}$. For every $i\in[m]$ we can write $[m]\setminus\{i\}=I\sqcup J$, where $|I|=i+1$ and $|J|=m-i-2$, hence $I\sqcup J^{\prime}$ is a maximal simplex in $\mathrm{Bier}(K)$, so $\mathrm{Bier}(K)$ is not a suspension with a vertex pair $\{i,i^{\prime}\}$ and therefore it is not a suspension. Finally, all the three chromatic numbers $\chi(K)$, $\chi(K^{\vee})$, and $\chi(\mathrm{Bier}(K))$ are equal to $m$, since the complete graph on $m$ vertices is a subcomplex of all the three complexes.
The next result is formally a generalization of Proposition 2.9 (b) and (c). However, in fact two statements are equivalent, since chromatic number of a simplicial complex depends only on its 1-skeleton.
@@
Note that one can add one color to any coloring of $K$ and color an apex (vertex pair) of the weak cone (weak suspension) $L$ over $K$ in that new color. This shows that
$\chi(L)\leq\chi(K)+1.$
On the other hand, for any coloring $C$ of $L$, the apex (vertex pair) of the weak cone (weak suspension) $L$ over $K$ is colored in colors not belonging to the restriction $C|_{K}$. This implies that
$\chi(K)<\chi(L).$
Combining the two inequalities above yields $\chi(L)=\chi(K)+1$, which finishes the proof. ∎
Lemma 3.8.
Let $K$ be a simplicial complex on $[m]$. Suppose $\mathrm{Bier}(K)$ is a weak suspension with a vertex pair $\{a,b\}$. Then there exists an $i\in[m]$ such that
$\{a,b\}=\{i,i^{\prime}\}.$
Proof.
Denote by $\Gamma=\mathrm{sk}^{1}(\mathrm{Bier}(K))$. Since $a$ and $b$ are not linked by an edge in $\mathrm{Bier}(K)$, one and only one of the following three possibilities takes place:
  • $\{a,b\}\in\mathrm{MF}(K)$;
  • $\{a,b\}\in\mathrm{MF}(K^{\vee})$;
  • $\{a,b\}=\{i,i^{\prime}\}$ for some $i,1\leq i\leq m$.
However, in the first case, one has:
$\phi(a)\in\operatorname{Link}_{\Gamma}(b)\text{ and }\phi(a)\notin\operatorname{Link}_{\Gamma}(a)$
and similarly in the second case, one has:
$\phi^{-1}(a)\in\operatorname{Link}_{\Gamma}(b)\text{ and }\phi^{-1}(a)\notin\operatorname{Link}_{\Gamma}(a).$
This shows that only the last possibility takes place. ∎
@@
If $\mathrm{Bier}(K)$ is a weak suspension with a vertex pair $\{i,i^{\prime}\}$, then $K$ and $K^{\vee}$ are weak cones with apex $i$ and $i^{\prime}$, respectively, since they are subcomplexes. On the other hand, if $K$ and $K^{\vee}$ are weak cones with apex $i$ and $i^{\prime}$, respectively, for all non-ghost vertices $j$ and $j^{\prime}$, $j\neq i$, all simplices $\{i,j\}$ and $\{i^{\prime},j^{\prime}\}$ are in $\mathrm{Bier}(K)$ and, by the construction of the Bier sphere also all simplices $\{i,j^{\prime}\}$ and $\{i^{\prime},j\}$. ∎
@@
The implication (a) $\Rightarrow$ (c) follows from [22, Lemma 3.3 (c)], where it was shown that
@@
The example in the following figure is the motivation for the next lemma.
222233331111
@@
Figure 5. On the left is the graph $\Gamma_{6}$, which is self-dual, and its Bier sphere, which is the $6$-cycle. On the right is a cone over $\Gamma_{6}$, which is also self-dual, and its Bier sphere, which is a suspension over $\mathrm{Bier}(\Gamma_{6})$.
@@
The implication $\Rightarrow$ follows directly from [22, Lemma 3.3 (c)]:
@@
Let us prove the implication $\Leftarrow$. Suppose $I\in\mathrm{M}(K)$ and $m\notin I$. Then $I\in\mathrm{M}(L)$, since $L$ is the deletion of $m$ from $K$. Therefore $m^{\prime}\in[m^{\prime}]\setminus I^{\prime}\in\mathrm{MF}(\mathrm{Bier}(K))$ and hence
$[(m-1)^{\prime}]\setminus I^{\prime}\notin\mathrm{MF}(\mathrm{Bier}(K)_{[m-1]\sqcup[(m-1)^{\prime}]}),$
@@
$[(m-1)^{\prime}]\setminus I^{\prime}\in\mathrm{MF}(\mathrm{Bier}(L)).$
Thus, $\mathrm{Bier}(L)$ and $\mathrm{Bier}(K)_{[m-1]\sqcup[(m-1)^{\prime}]}$ are different simplicial complexes on $[m-1]\sqcup[(m-1)^{\prime}]$, which contradicts our assumptions. This finishes the proof. ∎
@@
Suppose $\chi(\mathrm{Bier}(K))=m-1$ and let $c\colon V(\mathrm{Bier}(K))\to C=\{c_{1},c_{2},\ldots,c_{m-1}\}$ be a proper coloring. Without loss of generality, we can assume $f_{0}(K)<m$ and $\{m\}\notin K$. Therefore $[(m-1)^{\prime}]\in K^{\vee}$ and $C=\{c(1^{\prime}),\ldots,c((m-1)^{\prime})\}$. By Proposition 3.1 it follows that $\chi(K^{\vee})=m-1$.
@@
Thus, we conclude that $f_{0}(K)\leq m-2$, and there exists $j\in\{f_{0}(K)+1,\ldots,m-1\}$ such that $\{j^{\prime},m^{\prime}\}\not\in K^{\vee}$, since $\chi(K^{\vee})=m-1$. Furthermore, the vertex $j^{\prime}$ must be unique in this case, since having more than one such $j^{\prime}$ would imply the existence of two different $(m-3)$-dimensional simplices in $K$, which contradicts the condition $f_{0}(K)\leq m-2$.
Thus, we can assume that $\{(m-1)^{\prime},m^{\prime}\}\notin K^{\vee}$ and $\{j^{\prime},m^{\prime}\}\in K^{\vee}$ for all $21\leq j\leq m-$ 2. Hence finally $f_{0}(K)=m-2$ and moreover $K=\Delta_{[m-2]}$, with two ghost vertices $m-1$ and $m$. To complete the proper $(m-1)$-coloring $C$ of $\mathrm{Bier}(K)$ we set
@@
If $\{m^{\prime}\}\notin K^{\vee}$, then $K^{\vee}=\Delta_{[(m-1)^{\prime}]}$ and hence $K=\Delta_{[m-1]}$. In this case, obviously, $\chi(\mathrm{Bier}(K))=m-1$ with coloring $C$ given by:
@@
This final statement follows from the definition of a Bier sphere, completing the proof of the first part. The second part follows from Example 2.11. ∎
@@
Suppose $\mathrm{Bier}(K)$ is not a weak suspension and $\chi(\mathrm{Bier}(K))=m-1$. Let $c\colon V(\mathrm{Bier}(K))=[m]\sqcup[m^{\prime}]\to C$ be a proper coloring of $\mathrm{Bier}(K)$, where $|C|=m-1$.
@@
Since both $\{1,j^{\prime}\},\{2,j^{\prime}\}\in\mathrm{Bier}(K)$ for all $j\in\{3,\ldots,m\}$ it implies that $c(1),c(2)\notin C^{\prime}$ and therefore $c(1)=c(2)=:\alpha$, since $|C^{\prime}|=m-2$ and $|C|=m-1$. Thus, we have $C=C^{\prime}\sqcup\{\alpha\}$.
Suppose either $1$ or $2$ is not an isolated vertex of $K$. Without loss of generality, we can assume that $\{1,3\}\in K$. It implies that $c(3)\in C^{\prime}$. As $\{3,j^{\prime}\}\in\mathrm{Bier}(K)$ for all $j\in\{4,\ldots,m\}$ we get that $c(3)=c(3^{\prime})$.
Observe that $\{1,2^{\prime}\},\{1^{\prime},2\}\in\mathrm{Bier}(K)$ and hence $c(1^{\prime}),c(2^{\prime})\neq\alpha$. It follows that there are vertices $i^{\prime},j^{\prime}\in\{3^{\prime},4^{\prime},\ldots,m^{\prime}\}\in K^{\vee}$ such that $c(1^{\prime})=c(i^{\prime})$ and $c(2^{\prime})=c(j^{\prime})$. Hence the corresponding 2-element sets are non-edges in $K^{\vee}$ and Alexander duality yields:
$[m]\setminus\{1,i\},[m]\setminus\{2,j\}\in K.$
If $i\neq j$ the latter inclusions imply that $\{2,j\},\{1,i\}\in K$ and so $c(i),c(j)\neq\alpha$. Therefore
@@
$[m]\setminus\{1,i\},[m]\setminus\{2,i\},\text{ and }[m]\setminus\{1,2\}\in K,$
@@
If $c(i)\neq\alpha$, then $c(i)\in C^{\prime}$ and so $c(i)=c(i^{\prime})$ by definition of Bier sphere. Then $\{1^{\prime},i\}\in\mathrm{Bier}(K)$ and $c(1^{\prime})=c(i^{\prime})=c(i)$, see above, a contradiction. Thus, $c(i)=\alpha$.
This implies that $\{1,i\},\{2,i\}\notin K$ and given $x$ such that $3\leq x\neq i\leq m$ we have
@@
$c(1^{\prime})$ $=c(2^{\prime})=c(i^{\prime})=\beta\not\in C^{\prime},$
@@
However, in such a Bier sphere, the pairs $\{j,j^{\prime}\}$ for each $j\in[m]\setminus\{1,2,i\}$ can be taken as a weak suspension vertex pair and therefore $\mathrm{Bier}(K)$ is a weak suspension, which contradicts the assumptions of the statement.
@@
$[m^{\prime}]\setminus\{1^{\prime},i^{\prime}\},[m^{\prime}]\setminus\{2^{\prime},i^{\prime}\}\in K^{\vee}\text{ for each }i\in\{3,\ldots,m\}.$
Recall that we also have $[m^{\prime}]\setminus\{1^{\prime},2^{\prime}\}\in K^{\vee}$. Therefore, since $m\geq 4$ by the assumptions of the statement, we have $c(1^{\prime}),c(2^{\prime})\notin C^{\prime}=C\setminus\{\alpha\}$, which implies that $c(1^{\prime})=\alpha=c(2^{\prime})$, a contradiction as $\{1^{\prime},2\},\{1,2^{\prime}\}\in\mathrm{Bier}(K)$ and $\alpha=c(1)=c(2)$. Therefore, $\mathrm{Bier}(K)$ is a weak suspension, which finishes the proof. ∎
@@
Since $\mathrm{Bier}(K)$ is not a suspension, there exists a simplex $\sigma\in\operatorname{Del}_{K}(1)$ such that $\{1\}\cup\sigma\notin K$. Then, by definition of Alexander duality, its complement $[2^{\prime},3^{\prime},\ldots,m^{\prime}]\setminus\sigma^{\prime}\in K^{\vee}$ and therefore, by definition of Bier sphere, $\tau:=([2^{\prime},3^{\prime},\ldots,m^{\prime}]\setminus\sigma^{\prime})\cup\sigma$ is a maximal simplex in $\mathrm{Bier}(K)$.
By Proposition 3.9, $K$ is a weak cone with apex $1$, hence the vertex $1$ is linked by an edge in $\mathrm{Bier}(K)$ to each vertex of $\sigma\in K$ and therefore, to each vertex of $\tau$. Since $|\tau|=m-1$, it follows that the 1-skeleton of the full subcomplex of $\mathrm{Bier}(K)$ on the vertex set $\{1\}\cup\tau$ is a complete graph with $m$ vertices. This yields $\chi(\mathrm{Bier}(K))=m$, by Proposition 3.1, which finishes the proof. ∎
@@
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ and $m\geq 3$.
@@
To prove the implication $\Leftarrow$, we observe that by [22, Lemma 3.3 (c)] one has:
$\mathrm{Bier}(\Delta\ast\Gamma_{4})=\mathrm{Bier}(\Delta\ast G_{4})=\Sigma^{m-3}(Z_{4})\text{ and }\mathrm{Bier}(\Delta\ast\Gamma_{6})=\Sigma^{m-3}(Z_{6}).$
@@
$\gamma(\mathrm{Bier}(K))\leq(m-3)+\chi(P_{4})\text{ or }\gamma(\mathrm{Bier}(K))\leq(m-3)+\chi(P_{6}),$
respectively. Both right hand sides are equal to $m-1$. Application of Proposition 3.1 finishes the proof of this implication.
To prove the implication $\Rightarrow$, we argue by induction on $m$. For $m\leq 4$ the statement is true by the complete description of combinatorial Bier spheres in dimensions 1 and 2, see [21, Theorem 2.16]. Suppose the statement holds for a fixed $m\geq 4$, and let us show it is true when $K$ is a simplicial complex on $[m+1]$. Due to Lemma 3.12 it remains to consider the case when both $K$ and $K^{\vee}$ have no ghost vertices:
@@
By our assumption, $\chi(\mathrm{Bier}(K))=m$ and therefore, by Lemma 3.13, $\mathrm{Bier}(K)$ is a weak suspension. Then, by Lemma 3.14, $\mathrm{Bier}(K)$ is a suspension.

### v2

@@ 3. Bier spheres and their chromatic numbers
In this section, we show that every Bier sphere $\mathrm{Bier}(K)$ of a simplicial complex $K$ on $[m]$ with $m\geq 2$ has the chromatic number equal to either m1m-1, or $m$ and describe all the simplicial complexes $K$, for which $\chi(\mathrm{Bier}(K))=m-1$. The discussion of chromatic numbers of 1-dimensional Bier spheres in the end of the previous section as well as Example 2.11 will serve as the first step in the desired description.
@@
$i,i^{\prime}\mapsto c_{i}\text{ for }1\leq i\leq m,$
@@
Figure 3. The boundary of an octahedron is a suspension and therefore a weak suspension, but the boundaries of a tetrahedron and an icosahedron are not even weak suspensions.
Obviously, any cone over a simplicial complex is a weak cone complex, and any suspension over a simplicial complex is a weak suspension complex. The opposite implications fail; in the next example, we introduce a family of simplicial spheres, one sphere for each $m\geq 5$, such that each member of the family is a weak suspension, but not a suspension.
@@
By definition, its Alexander dual complex $K_{m}^{\vee}$ is a simplicial complex on the vertex set $m^{\prime}]$ given by
$\mathrm{M}(K_{m}^{\vee})=\{\{1^{\prime},m^{\prime}\},\{(m-1)^{\prime},m^{\prime}\},[(m-1)^{\prime}]\setminus\{i^{\prime}\}\,|\,2\leq i\leq m-1\}.$
@@
It is easy to see that $C(m)$ provides a proper coloring of both $K_{m}$ and $K_{m}^{\vee}$ with (m1)(m-1) colors, and so
@@
On the other hand, the 1-skeleta of the full subcomplexes $K_{m}\cap 2^{[m-1]}$ and $K^{\vee}_{m}\cap 2^{[(m-1)^{\prime}]}$ are complete graphs with (m1)(m-1) vertices, and therefore
@@
Skeleta of a simplex never yield a Bier sphere, which is a suspension, as shown in the following example. In the case of the 0-skeleton of the $m$-simplex and its Alexander dual, which is the (m3)(m-3)-skeleton of the $m$-simplex, the Bier sphere is not even a weak suspension, since it does not contain both edges $\{i,j\}$ and $\{i^{\prime},j^{\prime}\}$ for each pair of distinct vertices $i,j\in[m]$. The next example shows that in all other cases, the Bier sphere of a skeleton of a simplex is a weak suspension and not a suspension.
@@
Let $m\geq 5$, $1\leq k\leq m-4$, and $K=\Delta_{[m]}^{(k)}$, i.e., the set of maximal simplices is $\mathrm{M}(K)=\{I\subset[m]\,|\,|I|=k+1\}$. Then $I^{\prime}$ is a minimal non-face in $K^{\vee}$ if and only if |I|=mk1|I|=m-k-1, hence $K^{\vee}=\Delta_{[m^{\prime}]}^{(m-k-3)}$. For every $\in[m], iji\neq j$, we have $\{i,j^{\prime}\}$, $\{i^{\prime},j\}\in\mathrm{Bier}(K)$, since there are no ghost vertices in $K$ and $K^{\vee}$, $\{i,j\}\in K\subset\mathrm{Bier}(K)$, since $k\geq 1$, and $\{i^{\prime},j^{\prime}\}\in K^{\vee}\subset\mathrm{Bier}(K)$, since $m-k-3\geq 1$. Therefore $\mathrm{Bier}(K)$ is a weak suspension for every pair $\{i,i^{\prime}\}$. For every $i\in[m]$ we can write $m]\setminus\{i\}=I\sqcup J$, where $|I|=k+1$ and |J|=mk2|J|=m-k-2, hence $I\sqcup J^{\prime}$ is a maximal simplex in $\mathrm{Bier}(K)$, so $\mathrm{Bier}(K)$ is not a suspension with a vertex pair $\{i,i^{\prime}\}$ and therefore it is not a suspension. Finally, all three chromatic numbers $\chi(K)$, $\chi(K^{\vee})$, and $\chi(\mathrm{Bier}(K))$ are equal to $m$, since the complete graph on $m$ vertices is a subcomplex of all three complexes.
The next result is formally a generalization of Proposition 2.9 (b) and (c). However, in fact, the two statements are equivalent since the chromatic number of a simplicial complex depends only on its 1-skeleton, so we omit the proof.
@@
In the following lemma, we analyze vertex pairs of a Bier sphere being a weak suspension.
Lemma 3.8.
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 3$. Suppose $\mathrm{Bier}(K)$ is a weak suspension. Then there exists an $i\in[m]$ such that $\{i,i^{\prime}\}$ is a vertex pair of $\mathrm{Bier}(K)$.
@@
Denote by $\Gamma=\mathrm{sk}^{1}(\mathrm{Bier}(K))$. Assume $\{a,b\}\subseteq V(K)$ is a vertex pair of $\mathrm{Bier}(K)$. Then $\{a,b\}\in\mathrm{MF}(K)$ and $\{a^{\prime}\},\{b^{\prime}\}\in\mathrm{MF}(K^{\vee})$. Hence
[…]
The latter implies that $K^{\vee}$ is isomorphic to $\Delta_{(m-2)^{\prime}}$, since $\{a^{\prime}\},\{b^{\prime}\}\in\mathrm{MF}(K^{\vee})$. Then Example 2.11 shows the statement is true in this case. The case where $\{a^{\prime},b^{\prime}\}\subseteq V(K^{\vee})$ is a vertex pair of $\mathrm{Bier}(K)$ is similar. ∎
@@
If $\mathrm{Bier}(K)$ is a weak suspension with a vertex pair $\{i,i^{\prime}\}$, then $K$ and $K^{\vee}$ are weak cones with apex $i$ and $i^{\prime}$, respectively, since they are subcomplexes. On the other hand, if $K$ and $K^{\vee}$ are weak cones with apex $i$ and $i^{\prime}$, respectively, then for any geometrical vertices $j$ and $k^{\prime}$ of $\mathrm{Bier}(K)$ with $j,k\neq i$, the edges $\{i,j\}$ and $\{i^{\prime},k^{\prime}\}$ are in $\mathrm{Bier}(K)$ and, by the Bier sphere construction, also the edges $\{i,k^{\prime}\}$ and $\{i^{\prime},j\}$. ∎
@@
The implication (a) $\Rightarrow$ (c) follows from [28, Lemma 3.3 (c)], where it was shown that
@@
The example in the following figure serves as a motivation for the next lemma.
223311
@@
Figure 5. On the left is the graph $\Gamma_{6}$, which is self-dual, and its Bier sphere, which is the 66-cycle. On the right is a cone over $\Gamma_{6}$, which is also self-dual, and its Bier sphere, which is a suspension over $\mathrm{Bier}(\Gamma_{6})$.
@@
The implication $\Rightarrow$ follows directly from [28, Lemma 3.3 (c)]:
@@
Let us prove the implication $\Leftarrow$. Suppose $I\in\mathrm{M}(K)$ and $m\notin I$. Then $I\in\mathrm{M}(L)$, since $L$ is the deletion of $m$ from $K$. Therefore, $m^{\prime}\in[m^{\prime}]\setminus I^{\prime}\in\mathrm{MF}(\mathrm{Bier}(K))$ and hence
$(m-1)^{\prime}]\setminus I^{\prime}\notin\mathrm{MF}(\mathrm{Bier}(K)_{[m-1]\sqcup[(m-1)^{\prime}]}),$
@@
$(m-1)^{\prime}]\setminus I^{\prime}\in\mathrm{MF}(\mathrm{Bier}(L)).$
Thus, $\mathrm{Bier}(L)$ and $\mathrm{Bier}(K)_{[m-1]\sqcup[(m-1)^{\prime}]}$ are different simplicial complexes on $m-1]\sqcup[(m-1)^{\prime}]$, which contradicts our assumptions. This finishes the proof. ∎
@@
Suppose $\chi(\mathrm{Bier}(K))=m-1$ and let $c\colon V(\mathrm{Bier}(K))\to C=\{c_{1},c_{2},\ldots,c_{m-1}\}$ be a proper coloring. Without loss of generality, we can assume $f_{0}(K)<m$ and $\{m\}\notin K$. Therefore, $(m-1)^{\prime}]\in K^{\vee}$ and $C=\{c(1^{\prime}),\ldots,c((m-1)^{\prime})\}$. By Proposition 3.1, it follows that $\chi(K^{\vee})=m-1$.
@@
Thus, we conclude that $f_{0}(K)\leq m-2$, and there exists $j\in\{f_{0}(K)+1,\ldots,m-1\}$ such that $\{j^{\prime},m^{\prime}\}\not\in K^{\vee}$, since $\chi(K^{\vee})=m-1$. Furthermore, the vertex $j^{\prime}$ must be unique in this case, since having more than one such $j^{\prime}$ would imply the existence of two different (m3)(m-3)-dimensional simplices in $K$, which contradicts the condition $f_{0}(K)\leq m-2$.
Thus, we can assume that $\{(m-1)^{\prime},m^{\prime}\}\notin K^{\vee}$ and $\{j^{\prime},m^{\prime}\}\in K^{\vee}$ for all $1\leq j\leq m-2$. Hence finally $f_{0}(K)=m-2$ and moreover $K=\Delta_{[m-2]}$, with two ghost vertices, m1m-1 and $m$. To complete the proper (m1)(m-1)-coloring $c$ of $\mathrm{Bier}(K)$ we set:
@@
If $\{m^{\prime}\}\notin K^{\vee}$, then $K^{\vee}=\Delta_{[(m-1)^{\prime}]}$ and hence $K=\Delta_{[m-1]}$. In this case, obviously, $\chi(\mathrm{Bier}(K))=m-1$ with the proper (m1)(m-1)-coloring $c$ given by:
@@
This final statement follows from the definition of a Bier sphere, completing the proof of the necessity implication. The sufficiency implication follows from Example 2.11. ∎
@@
Suppose $\mathrm{Bier}(K)$ is not a weak suspension and $\chi(\mathrm{Bier}(K))=m-1$. Let $c\colon V(\mathrm{Bier}(K))=[m]\sqcup[m^{\prime}]\to C$ be a proper coloring of $\mathrm{Bier}(K)$, where |C|=m1|C|=m-1.
@@
Since both $\{1,j^{\prime}\},\{2,j^{\prime}\}\in\mathrm{Bier}(K)$ for all $j\in\{3,\ldots,m\}$, it implies that $c(1),c(2)\notin C^{\prime}$ and therefore $c(1)=c(2)=:\alpha$, since $|C^{\prime}|=m-2$ and |C|=m1|C|=m-1. Thus, we have $C=C^{\prime}\sqcup\{\alpha\}$.
Suppose either $1$ or $2$ is not an isolated vertex of $K$. Without loss of generality, we can assume that $\{1,3\}\in K$. It implies that $c(3)\in C^{\prime}$. As $\{3,j^{\prime}\}\in\mathrm{Bier}(K)$ for all $j\in\{4,\ldots,m\}$, we get that $c(3)=c(3^{\prime})$.
Observe that $\{1,2^{\prime}\},\{1^{\prime},2\}\in\mathrm{Bier}(K)$ and hence $c(1^{\prime}),c(2^{\prime})\neq\alpha$. It follows that there are vertices $i^{\prime},j^{\prime}\in\{3^{\prime},4^{\prime},\ldots,m^{\prime}\}\in K^{\vee}$ such that $c(1^{\prime})=c(i^{\prime})$ and $c(2^{\prime})=c(j^{\prime})$. Hence, the corresponding 2-element sets are non-edges in $K^{\vee}$ and Alexander duality yields:
$m]\setminus\{1,i\},[m]\setminus\{2,j\}\in K.$
If $i\neq j$ the above inclusions imply that $\{2,j\},\{1,i\}\in K$, and so $c(i),c(j)\neq\alpha$. Therefore,
@@
$m]\setminus\{1,i\},[m]\setminus\{2,i\},\text{ and }[m]\setminus\{1,2\}\in K,$
@@
If $c(i)\neq\alpha$, then $c(i)\in C^{\prime}$ and so $c(i)=c(i^{\prime})$ by definition of a Bier sphere. Then $\{1^{\prime},i\}\in\mathrm{Bier}(K)$ and $c(1^{\prime})=c(i^{\prime})=c(i)$, see above, a contradiction. Thus, $c(i)=\alpha$.
This implies that $\{1,i\},\{2,i\}\notin K$ and given $x$ such that $3\leq x\neq i\leq m$, we have:
@@
$c(1^{\prime})$ $=c(2^{\prime})=c(i^{\prime})=\beta\in C^{\prime},$
@@
However, in such a Bier sphere, the pairs $\{j,j^{\prime}\}$ for each $j\in[m]\setminus\{1,2,i\}$ can be taken as a weak suspension vertex pair and therefore $\mathrm{Bier}(K)$ is a weak suspension, which contradicts our initial assumption.
@@
$m^{\prime}]\setminus\{1^{\prime},i^{\prime}\},[m^{\prime}]\setminus\{2^{\prime},i^{\prime}\}\in K^{\vee}\text{ for each }i\in\{3,\ldots,m\}.$
Recall that we also have $m^{\prime}]\setminus\{1^{\prime},2^{\prime}\}\in K^{\vee}$. Therefore, since $m\geq 4$ by the assumptions of the statement, we have $c(1^{\prime}),c(2^{\prime})\notin C^{\prime}=C\setminus\{\alpha\}$, which implies that $c(1^{\prime})=\alpha=c(2^{\prime})$, a contradiction, as $\{1^{\prime},2\},\{1,2^{\prime}\}\in\mathrm{Bier}(K)$ and $\alpha=c(1)=c(2)$. Thus, $\mathrm{Bier}(K)$ is a weak suspension, which finishes the proof. ∎
@@
Since $\mathrm{Bier}(K)$ is not a suspension, there exists a simplex $\sigma\in\operatorname{Del}_{K}(1)$ such that $\{1\}\cup\sigma\notin K$. Then, by definition of Alexander duality, its complement $2^{\prime},3^{\prime},\ldots,m^{\prime}]\setminus\sigma^{\prime}\in K^{\vee}$ and therefore, by definition of a Bier sphere, $\tau=([2^{\prime},3^{\prime},\ldots,m^{\prime}]\setminus\sigma^{\prime})\cup\sigma$ is a maximal simplex in $\mathrm{Bier}(K)$.
By Proposition 3.9, $K$ is a weak cone with apex $1$. Hence the vertex $1$ is linked by an edge in $\mathrm{Bier}(K)$ to each vertex of $\sigma\in K$ and, therefore, to each vertex of $\tau$. Since $|\tau|=m-1$, it follows that the 1-skeleton of the full subcomplex of $\mathrm{Bier}(K)$ on the vertex set $\{1\}\cup\tau$ is a complete graph with $m$ vertices. This yields $\chi(\mathrm{Bier}(K))=m$, by Proposition 3.1, which finishes the proof. ∎
@@
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 3$.
@@
To prove the implication $\Leftarrow$, we observe that by [28, Lemma 3.3 (c)] one has:
$\mathrm{Bier}(\Delta^{m-4}\ast\Gamma_{4})=\mathrm{Bier}(\Delta^{m-4}\ast G_{4})=\Sigma^{m-3}(Z_{4})\text{ and }\mathrm{Bier}(\Delta^{m-4}\ast\Gamma_{6})=\Sigma^{m-3}(Z_{6}).$
@@
$\chi(\mathrm{Bier}(K))\leq(m-3)+\chi(P_{4})\text{ or }\chi(\mathrm{Bier}(K))\leq(m-3)+\chi(P_{6}),$
respectively. Both right hand sides are equal to m1m-1. Application of Proposition 3.1 finishes the proof of this implication.
To prove the implication $\Rightarrow$, we argue by induction on $m$. For $m\leq 4$ the statement is true by the complete description of combinatorial Bier spheres in dimensions 1 and 2, see Example 2.3 and [27, Theorem 2.16]. Suppose the statement holds for a fixed $m\geq 4$, and let us show it is true when $K$ is a simplicial complex on $[m+1]$ and $\chi(\mathrm{Bier}(K))=m$. Due to Lemma 3.12, it remains to consider the case when both $K$ and $K^{\vee}$ have no ghost vertices:
@@
Therefore, by Lemma 3.13, $\mathrm{Bier}(K)$ is a weak suspension. Then, by Lemma 3.14, $\mathrm{Bier}(K)$ is a suspension.